# Holographic spectral function of fermion in instantonic plasma

### v1

@@ 3.2 Covariant Dirac equation for a probe fermion
Keep the formulas in Section 3.1 in hand, let us investigate a probe fermion in the 5d holographic background (2.9). In general, the 5d homogeneous background metric can be written as,
@@
where $g_{tt},g_{xx},g_{zz}$ depend on $z$ only. For a probe fermion in bulk, its dynamic is described by the Dirac action given as,
$S_{\mathrm{bulk}}=i\int d^{5}x\sqrt{-g}\bar{\psi}\left(\Gamma^{M}\nabla_{M}-m\right)\psi,$ (3.7)
so the Dirac equation reads
@@
$\Gamma^{M}\nabla_{M}=\frac{1}{\sqrt{-g_{tt}}}\gamma^{0}\partial_{0}+\frac{1}{\sqrt{g_{xx}}}\gamma^{i}\partial_{i}+\frac{1}{\sqrt{g_{zz}}}\gamma\partial_{z}+\frac{1}{4g_{xx}g_{tt}\sqrt{g_{zz}}}\left[3g_{tt}g_{xx}^{\prime}+g_{xx}g_{tt}^{\prime}\right]\gamma.$ (3.10)
@@ 3.3 Construction for the correlation function
Figure 1: The real and imaginary parts of zero-th order correlation functions $G_{R,\alpha}^{\left(0\right)}$ as functions of $\omega,\mathrm{k}$. The parameter is chosen as $q=0,m=0.01,z_{H}=1.$
Figure 2: Density plots of the zero-th order correlation functions $G_{R,\alpha}^{\left(0\right)}$ with the parameters as $q=0,m=0.01,z_{H}=1.$ The white regions refer to the peaks in the correlation functions.
@@
Therefore the correlation function could be evaluated numerically by solving (3.33) with the infalling boundary condition $G_{\alpha}|_{z=z_{H}}=\left(-1\right)^{\alpha}i$ obtained by analyzing its asymptotic behavior near $z=z_{H}$ as [35, 36, 37, 38, 39]. Furthermore, when the D-instanton is taken into account, the correlation function given in (3.32) should also depend on the charge density of the D-instanton denoted as $q$. In order to evaluate exactly the $q$-dependence in the correlation function, we consider its leading order correction by expand $G_{\alpha}$ as,
@@
And $\xi_{\alpha}$ is the leading order correction satisfying the equations
@@
By analyzing the asymptotic behavior of (3.36), we find the infalling boundary condition for $\mathcal{G}_{\alpha}$ could be $\mathcal{G}_{\alpha}|_{z=z_{H}}=\left(-1\right)^{\alpha}i$ which is exactly same as the infalling boundary condition for $G_{\alpha}$. Therefore, the boundary condition for $\xi_{\alpha}$ is expected as $\xi_{\alpha}|_{z=z_{H}}=0$. Altogether, we will solve (3.36) and (3.37) numerically in order to evaluate the associated correlation function $G_{R}$ as the fermionic spectral function in holography.
4 The numerical analysis
Figure 3: The real and imaginary parts of leading order correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega,\mathrm{k}$ with $m=0.01,z_{H}=1.$
Figure 4: Density plots of leading order correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega,\mathrm{k}$ with $m=0.01,z_{H}=1.$ The white regions refer to the peaks in the correlation functions
Figure 5: The real and imaginary parts of leading order correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega$ with various $\mathrm{k}$ and $m=0.01,z_{H}=1$. There are always two peaks for various $\mathrm{k}$ in the correlation function which correspond to the two branches of the dispersion curves.
We write down the correlation functions presented in (3.32) as the combinations of its zero-th order $G_{R,\alpha}^{\left(0\right)}$ and first order solution $G_{R,\alpha}^{\left(1\right)}$ as,
@@ 5 Summary
In this work, we study the correlation function of fermion in the D(-1)-D3 brane system which describes holographically the hot plasma with instantonic configuration. To simplify the setup, we briefly reduce the 10d supergravity background to an 5d background of gravity-dilaton-axion system. By probing a spinor in the 5d bulk, we derive the equations for the holographic correlation function from the Dirac equation, then demonstrate the numerical calculations with the infalling boundary condition. In particular, we expand the correlation function as zero-th order solution plus first order solution. Our numerical results illustrate that the first order correction to the total correlation function including two branches of the fermionic dispersion curves is dominated for small momentum, while the zero-th order correlation function dominates the behavior of the fermionic dispersion curves at large momentum giving only one merged dispersion curves on light cone. This behavior is out of reach without the presence of the instanton and it is very close to the fermionic dispersion curves obtained by HTL. Furthermore, we also find the effective mass generated by the medium effect of fermion splits into two values due to the spin-dependent interactions among fermions induced by instantons which has be studied in the framework of QCD with instanton [1, 2]. Altogether, it means seemingly the instantonic configuration in the plasma is very influential for the spectral function of the plasmino.
@@ Acknowledgements
We would like to thank Prof.Lin Shu for helpful discussion. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12005033 and the Fundamental Research Funds for the Central Universities under Grant No. 3132024192.

### v2

@@ 4.2 Covariant Dirac equation for a probe fermion
Keep the formulas in Section 3.1 in hand, let us investigate a probe fermion in the D+1-dimensional holographic background (2.10). In general, the D+1-dimensional homogeneous background metric can be written as,
@@
where $g_{tt},g_{xx},g_{zz}$ depend on $z$ only. For a probe fermion in bulk, its dynamic is described by the Dirac action given in (3.2), so the Dirac equation reads
@@
$\Gamma^{M}\nabla_{M}=\frac{1}{\sqrt{-g_{tt}}}\gamma^{0}\partial_{0}+\frac{1}{\sqrt{g_{xx}}}\gamma^{i}\partial_{i}+\frac{1}{\sqrt{g_{zz}}}\gamma\partial_{z}+\frac{1}{4g_{xx}g_{tt}\sqrt{g_{zz}}}\left[\left(D-1\right)g_{tt}g_{xx}^{\prime}+g_{xx}g_{tt}^{\prime}\right]\gamma.$ (4.9)
@@ 4.3 Construction for the correlation function
Figure 2: The real and imaginary parts of zero-th order correlation functions $G_{R,\alpha}^{\left(0\right)}$ as functions of $\omega,\mathrm{k}$ in the deconfined phase. The parameter is chosen as $q=0,m=0.01,z_{H}=1.$
Figure 3: Density plots of the zero-th order correlation functions $G_{R,\alpha}^{\left(0\right)}$ in the deconfined phase with the parameters as $q=0,m=0.01,z_{H}=1.$ The white regions refer to the peaks in the correlation functions.
@@
Therefore the correlation function could be evaluated numerically by solving (4.32) with the infalling boundary condition $G_{\alpha}|_{z=z_{H}}=\left(-1\right)^{\alpha}i$ obtained by analyzing its asymptotic behavior near $z=z_{H}$ as [35, 36, 39, 40, 41].
4.4 The numerical analysis
Deconfined phase
When the D-instanton is taken into account, the correlation function given in (4.31) should also depend on the charge density of the D-instanton denoted as $q$. In order to evaluate exactly the $q$-dependence in the correlation function, for $q\ll 1$ we consider its leading order correction by expand $G_{\alpha}$ as,
@@ 4.4 The numerical analysis
Here we have imposed the deconfined geometry (2.10) into (4.32), and $\xi_{\alpha}$ is the leading order correction satisfying the equations
@@
By analyzing the asymptotic behavior of (4.35), we find the infalling boundary condition for $\mathcal{G}_{\alpha}$ could be $\mathcal{G}_{\alpha}|_{z=z_{H}}=\left(-1\right)^{\alpha}i$ which is exactly same as the infalling boundary condition for $G_{\alpha}$. Therefore, the boundary condition for $\xi_{\alpha}$ is expected as $\xi_{\alpha}|_{z=z_{H}}=0$. Altogether, we will solve (4.35) and (4.36) numerically in order to evaluate the associated correlation function $G_{R}$ as the fermionic spectral function in holography. Write down the correlation functions presented in (4.31) as the combinations of its zero-th order $G_{R,\alpha}^{\left(0\right)}$ and first order solution $G_{R,\alpha}^{\left(1\right)}$ as,
@@
Figure 4: The real and imaginary parts of leading order deconfined correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega,\mathrm{k}$ with $m=0.01,z_{H}=1.$
Figure 5: Density plots of leading order deconfined correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega,\mathrm{k}$ with $m=0.01,z_{H}=1.$ The white regions refer to the peaks in the correlation functions
Figure 6: The real and imaginary parts of leading order deconfined correlation functions $G_{R,\alpha}^{\left(1\right)}$ as functions of $\omega$ with various $\mathrm{k}$ and $m=0.01,z_{H}=1$. There are always two peaks for various $\mathrm{k}$ in the correlation function which correspond to the two branches of the dispersion curves.
@@
Confined phase
[…]
Figure 7: The density plot of the holographic correlation function in confined phase with $q=1$. The white region refers to the peaks in the correlation function.
Figure 8: The holographic correlation function as a function of $\omega$ at $\mathrm{k}=0$. in confined phase with $q=1$.
@@ 5 Summary
In this work, we study the spectroscopy of baryonic fermion in the D(-1)-D3 brane system. The geometric background of this system given by IIB supergravity includes a black and a bubble solution which can describe respectively a deconfined and a confined gauge theory with instantons in holography. To simplify the setup, we briefly reduce the 10d supergravity background to an 5d background of gravity-dilaton-axion system. By probing a spinor in the 5d bulk, we illustrate the decomposition and dimensional reduction of the spinor, then obtain the fermionic spectrum in the confined geometry. Since the fermion must be a gauge-invariant operator according to the gauge-gravity duality, we further identify the fermion to baryon in the confined geometry and compare our mass spectrum with the baryonic experimental data. However, as most works about holography [40, 41], the gauge-invariant fermion in the deconfined phase is usually identified to baryonic plasmino instead of baryon since quarks and gluons can not form a confined baryon. Afterwards, we derive the equations for the holographic correlation function from the Dirac equation, then demonstrate the numerical calculations with the infalling boundary condition both in the deconfined and confined background. In particular, we expand the correlation function as zero-th order solution plus first order solution in the deconfined geometry. Our numerical results in the deconfined case illustrate that the first order correction to the total correlation function including two branches of the fermionic dispersion curves is dominated for small momentum, while the zero-th order correlation function dominates the behavior of the fermionic dispersion curves at large momentum giving only one merged dispersion curves on light cone. This behavior is out of reach without the presence of the instanton and it is very close to the fermionic dispersion curves obtained by HTL. Furthermore, we also find the effective mass generated by the medium effect of fermion splits into two values which may be due to the spin-dependent interactions among fermions induced by instantons. Although it has be discussed qualitatively in the framework of QCD with instanton [1, 2], it is interesting the holographic approach reproduces this result. Besides, in the confined case, the holographic correlation function explicitly indicates that the position of the onset mass is consistent with the fermionic spectrum. Altogether, it means seemingly the instantonic configuration in QCD is very influential for the spectroscopy of the fermion.
@@ Acknowledgements
We would like to thank Prof.Lin Shu for helpful discussion. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12005033 and the Fundamental Research Funds for the Central Universities under Grant No. 3132025192.